OpenStudy (anonymous):

Calculate the change in internal energy of 2kg of water at 90 degree celcius when it is changed to 3:30m3 of steam at 100 degrees. The whole process occurs at atmospheric pressure. The latent heat of vaporization of water is 2:26 x 10^6 J/kg.

OpenStudy (anonymous):

maybe 4604000 J just did it using some of my own values :p

OpenStudy (mos1635):

please elaborate...

OpenStudy (anonymous):

m*c*t=Q 2*4200*10=84000 m*l=Q 2*2.26x10^6=4520000 Total= 4520000+ 84000 4604000J

OpenStudy (anonymous):

argh i dont know if has to do something with the Charles and boyles laws

OpenStudy (mos1635):

that is exactly my thinking but that 3:30m3 is confusing.....

OpenStudy (mos1635):

f 2Kgr of Water terned in gas in P=1 atm T=100+273 R=0.082 n=1000/18 should have volule V=1.699 m3

OpenStudy (anonymous):

n=2000/18

OpenStudy (mos1635):

oooooppsssss

OpenStudy (mos1635):

and i did say 2 Kgr of water!!!! LOL

OpenStudy (mos1635):

4604000J it is then...

OpenStudy (anonymous):

pv=nRT is ideal gas equation i doubt if water ( steam) is an ideal gas ;)

OpenStudy (mos1635):

in higt temp and low densidy that is an acceptable aproximation

OpenStudy (mos1635):

aply PV=nRT gives 3.39844444 m3 close enough

OpenStudy (anonymous):

@mos1635 @hashsam1 hello guys why are u relating the solution to PV = nRt and how do u arrive at it values

OpenStudy (mos1635):

i was just cheking if given value of volume had any significance to the question. No need to conclude that in your answer.

OpenStudy (mos1635):

if we ignore given volume we have 1st water from 90 to water 100 Q1=m*c*ΔΤ1 m=2 Kgr c=4200 joule/Kgr*K ΔT1=10 2nd water from 100 to steam 100 Q2=m*l m= 2Kgr l=2.26 x 10^6 J/kg.

OpenStudy (anonymous):

thanks @mos1635